# Age of Job Completion Minimization with Stable Queues

### v1

@@ I Introduction
Due to the randomness in the state information of the edge computing device, we can model it as a stochastic process. The problem of tracking or remotely estimating such stochastic processes has been extensively studied in the literature. In particular, significant attention has been given to the remote estimation of Markov processes. Numerous studies have explored various aspects of this problem, including optimal sampling, estimation policies, and scheduling strategies, e.g.. see [maatouk2020age, kam2020age, chen2021scheduling, kriouile2022minimizing], and the references therein. In all these works, the sampling process does not change the dynamics of the underlying Markov chain. Thus, it is important to note that the aforementioned studies do not consider the presence of an external control that can actively influence the state evolution of the system. However, in edge computing scenarios, each job offloading event may alter the state of the device, thereby introducing a control-dependent dynamic. Motivated by this, in the literature, recent studies investigate the problem of tracking a Markov process under external control inputs [banerjee2025tracking, liyanaarachchi2025optimum, chamoun2025edge, liyanaarachchi2025age, sariisik2025maximize, chamoun2025mappo]. These studies are mainly done in the context of job offloading to computing devices whose dynamics can be modeled as a Markov process. Consequently, the term Markov machine has been introduced in the literature to describe such devices.
In this work, we consider a job-assignment system, where $N$ users present in the system stochastically submit their job requests to a central server, as shown in Fig. 1. For each user, the central server has a job queue, and it stores the job requests of a user in its dedicated job queue. The system includes a Markov machine that serves a job request assigned by the central server. At the beginning of a time slot, the server may sample the state information of the machine, paying a sampling cost, and if it observes that the machine is in a free state, it schedules a user and submits a job from the corresponding job queue to the machine. As we are considering a job-offloading system, our main goal is to maximize the number of jobs completed per unit time, and to this end, we introduce a novel metric termed the age of job completion. As will be discussed later, neither this metric nor a similar metric has been considered in the existing literature on job assignment to Markov machines. To minimize the age of job completion and the sampling cost, we propose two policy pairs, each comprising a scheduling policy and a sampling policy. For both policy pairs, we numerically find their performance. Furthermore, we derive sufficient conditions that ensure the stability of the job queues for both policy pairs. To the best of our knowledge, such stability results have not been explored in prior studies on job offloading to Markov machines.
@@ II Related Works
In [banerjee2025tracking], the authors considered a Markov machine equipped with a sampler that samples and transmits its state information to a central server at the cost of sampling. To quantify the tracking performance, the authors introduced a tailored version of the age of incorrect information (AoII) metric. The AoII metric does not directly address the maximization of the number of completed jobs, whereas the metric proposed here, i.e., the age of job completion metric, better captures this objective in job-offloading systems.
@@
In [liyanaarachchi2025optimum], the authors considered a network of Markov machines, where the jobs arrive at a resource allocator following a Poisson process. The authors studied an optimal sampling rate allocation problem for query-based sampling of the machines. To monitor the tracking performance, the authors used the well-known binary freshness (BF) metric. They also introduced the false rejection ratio (FRR) and the false acceptance ratio (FAR) metrics to quantify the effects of limited tracking: the former measures the fraction of jobs rejected despite the presence of a free machine due to insufficient tracking, while the latter measures the fraction of jobs accepted even though no machine was actually free, again due to limited tracking. It is important to note that, similar to AoII, none of these metrics directly captures the objective of maximizing the number of completed jobs. Moreover, the authors in [liyanaarachchi2025optimum] did not consider a job queue at the resource allocator. In [sariisik2025maximize], the authors considered a job assignment system with a network of exhausted workers and a resource allocator, where the efficiency of each worker depends on its current state. Similar to [liyanaarachchi2025optimum], the authors formulated and studied an optimal sampling rate allocation problem for query-based sampling of these workers, aiming to maximize the probability that each worker operates in its most efficient state.
@@
A revenue maximization problem has been considered in a job assignment system in [liyanaarachchi2025age]. They considered that a resource allocator has a unit-size buffer, i.e., it can store at most one job at a time. In [chamoun2025mappo], the authors considered a job-dispatching system with a network of dispatchers and edge servers; these servers were modeled as Markov machines. Their goal was to maximize the differential gain between job completion and sampling cost, and they formulated the problem as a discounted Markov decision process (MDP). As a result, their work did not directly maximize the average number of jobs completed over the time horizon. When a job was dispatched to a server, it was stored in a limited-size job queue, and if the queue was full, the oldest job was replaced. Similar to [chamoun2025mappo], the authors of [chamoun2025edge] considered a job-dispatching system for a network of dispatchers and edge servers; however, in contrast to [chamoun2025mappo], they did not consider a job queue at the edge server, i.e., if a dispatcher offloads a job to an edge server and if it is busy, the job is dropped. They proposed an index-based policy that maximizes the probability that a dispatched job is accepted by an edge server. However, in our work, the central server first samples the state of the machine and submits a job only if the machine is free, resulting in an acceptance probability of $1$ for offloaded jobs. Thus, our focus is on maximizing the number of jobs completed per unit time, rather than the job acceptance probability as in [chamoun2025edge].
@@
In contrast to [banerjee2025tracking, liyanaarachchi2025optimum, chamoun2025edge, liyanaarachchi2025age, sariisik2025maximize, chamoun2025mappo], in this work, we consider limitless full job queues, making the study of queue stability a necessary aspect of our problem.
@@ III System Model and Problem Formulation
In our job assignment system, user $i$ submits a job request to the central server at the end of each time slot, according to an i.i.d. Bernoulli random variable $a_{i}(t)$, with probability of success being $p_{i}>0$. When the machine is not executing a job assigned by the server, denoted as an external job, the machine can be either free or busy processing an internal job. We model the dynamics between these free and internally busy states with a binary symmetric Markov chain, with probability of state transition being q>0q>0.111For simplicity of computations, in this work, we consider a symmetric state transition, and the general case will be considered in a journal version.
@@
The machine only accepts an external job when it is in the free state. Thus, when the machine is not executing an external job and the server intends to assign a job to the machine, the server samples the state information of the machine, which incurs a cost $L$, and receives the sampled state information instantaneously. We denote this sampling operation at time $t$ with an indicator variable $\mu(t)$, where $\mu(t)=1$ implies that the server samples the state information of the machine at time $t$, and $\mu(t)=0$ implies otherwise. If the server infers that the machine is in a free state, the server instantaneously assigns a job to the machine from one of the non-empty job queues. At time $t$, we denote this scheduling operation with a vector $(\pi_{i}(t))_{i=1}^{N}$. At time $t$, if the server schedules a job from the job queue of user $i$, then $\pi_{i}(t)=1$; otherwise $\pi_{i}(t)=0$. Thus, we represent an employable policy by the central server as $\phi=(\mu(t),(\pi_{i}(t))_{i=1}^{N})_{t=1}^{\infty}$.
@@
We assume that the server performs the sampling, scheduling and job assignment processes, at the beginning of a time slot and the machine completes either an internal job or an external job at the end of a time slot. When the machine is processing a job of user $i$, the service time follows a geometric distribution with parameter $q_{i}>0$. During this period, the machine cannot process any other job, and the server is aware of the state of the machine. Once the job is completed, the server is instantaneously notified. After the machine completes a job, it transits to the internally busy state with probability $s>0$, or to the free state with probability (1s)(1-s). Thus, if an external job is completed at time slot t1t-1, the server becomes uncertain about the state of the machine at the beginning of slot $t$ and must sample its state information before submitting a future job. At time $t$, we denote the last known state information of the machine at the server with $x(t)\in\{-1,0,\ldots,N,f_{r}\}$. Here 1-1 corresponds to the internally busy state, the state 0 denotes the ambiguity of the state of the machine after it completes a job, $i\in\{1,\ldots,N\}$ indicates that the machine is processing a job of user $i$, and $f_{r}$ denotes that the machine is free. Note that, $x(t)$ gets updated if the server samples the state of the machine at time $t$, or receives a job completion acknowledgment at t1t-1, otherwise $x(t)$ evolves as x(t)=x(t1)x(t)=x(t-1). Note that state $f_{r}$ is only achieved when the sampled state of the machine is free and the server does not schedule any job. Furthermore, $x(t)$ remains 0 from the time a job is completed until the server next samples the state information. Similarly, with $x_{a}(t)$, we denote the actual state of the machine at time $t$. The variable $x_{a}(t)$ takes the values $\{-1,1,2\cdots,N,f_{r}\}$, retaining the same interpretation as $x(t)$. Fig. 2 shows the dynamics of the considered Markov machine.
@@
Intuitively, a policy $\phi$ that minimizes (3) also maximizes the number of completed jobs per unit time. This argument is formally shown in a subsequent work. In this work, first we present an adaptive randomized policy that comprises an adaptive randomized sampling policy and an adaptive randomized scheduling policy. Then, we study the max-age scheduling policy together with an adaptive randomized sampling policy. At time $t$, we observe the non-empty job queues and schedule the corresponding users with a stationary randomized policy. Whenever an identical set of job queues becomes non-empty, the same stationary randomized policy is applied. We call this policy an adaptive randomized scheduling policy. Similarly, based on the non-empty job queues, we define an adaptive randomized sampling policy. We provide a formal definition of an adaptive randomized policy in Definition 1.
@@
An adaptive stationary randomized policy can be viewed as a collection of pairs of stationary randomized policies - one governing user selection and one governing sampling - each supported on a distinct subset of $[N]$.
In a conventional stationary randomized scheduling policy, the central server schedules a user with a fixed probability throughout the entire time horizon, e.g., see [kadota18]. However, at a given time slot, a job queue can be empty; thus, scheduling that user during that slot with a non-zero probability would waste system resources. Thus, using system resources efficiently while being mathematically tractable motivates the proposed adaptive randomized policy. Similar to the definition of an adaptive randomized policy in Definition 1, we formally define an adaptive randomized sampling policy.
@@
$\bar{\Pi}=\{\,\mu(\mathcal{S}),:\mathcal{S}\subseteq[N],\,\mathcal{S}\neq\varnothing\,\},$ (8)
@@
We proceed by formally defining the notion of queue stability [tassiulas1990stability]. We denote $\bm{Q}(t)=\left(Q_{1}(t),\>Q_{2}(t),\ldots,Q_{N}(t)\right)$, and model the queue length process $\bm{Q}=\{\bm{Q}(t)\}_{t\geq 1}$ as a discrete-time Markov chain evolving over the state space $\mathcal{Q}=\mathbb{N}_{\geq 0}^{N}$.
@@
For an irreducible queue length process $\bm{Q}$, the system of queues is said to be stable if $\bm{Q}$ is positive recurrent.
Remark 2.
This notion of stability is equivalent to strong stability as defined in [neely2010stochastic], that is,
$\limsup_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{N}\mathbb{E}[Q_{i}(t)]<\infty.$ (9)
@@ IV-C Stabilizing the Queues
We now derive sufficient conditions for queue stability under both an adaptive randomized policy and a max-age policy. In both cases, the queue-length process evolves as a Markov chain and we follow Definition 3 for stability.
@@
stabilizes the queues if there exists $\epsilon>0$ such that, for every non-empty subset $\mathcal{S}\subseteq[N]$,
@@
Note that $1-\chi(q,s)$ provides a lower bound on the probability that the machine is free when sampled. With this interpretation, the conditions in Proposition 1 ensure that the total arrival rate does not exceed the worst-case effective service rate under any of the stationary randomized policies. Conditions of this nature are, in general, known to be necessary and sufficient for rate stability [neely2010stochastic]. Here, we further show that they also suffice for strong stability (see Remark 2) under the nontrivial dynamics of the considered model, implying positive recurrence of the underlying Markov chain and hence finite expected queue lengths in steady state.
@@
Under any adaptive randomized policy, the queues are stable if there exists $\epsilon>0$, such that
@@
Consider a system with $N$ users having arrival rates $p_{i}$, service rates $q_{i}$, and machine parameters $q$ and $s$. A max-age policy $\bar{\phi}=\left(\mu,\pi^{MA}\right)$ stabilizes the queues if there exists $\epsilon>0$ such that for every non-empty subset $\mathcal{S}\subseteq[N]$,
@@ V Numerical Results
We observe numerically that queue stability is achieved only when the users’ arrival rates are well below their respective service rates. For example, with model parameters $N=4$, $q=0.35$, $s=0.3$, $L=5$, service rates $\bar{\bm{q}}=[0.55\;0.73\;0.84\;0.91]$, and arrival rates $\bm{p}=[0.09\;0.09\;0.12\;0.14]$, the conditions in Propositions 1 and 2 are not satisfied and the queues are found numerically to not be stable. Furthermore, the conditions stated in Propositions 1 and 2 are observed to be only sufficient and not necessary for stability. For model parameters $N=4$, $q=0.5$, $s=0.5$, $L=5$, service rates $\bar{\bm{q}}=[0.4\;0.6\;0.8\;0.94]$ and arrival rates $\bm{p}=[0.04\;0.05\;0.06\;0.06]$, the sufficient stability conditions are not satisfied; however, numerical results show that the queues remain stable both under $\phi_{1}$ and $\bar{\phi}_{1}$.

### v2

@@ I Introduction
Due to the randomness in the state information of the edge computing device, we can model it as a stochastic process. The problem of tracking or remotely estimating such stochastic processes has been extensively studied in the literature. In particular, significant attention has been given to the remote estimation of Markov processes. Numerous studies have explored various aspects of this problem, including optimal sampling, estimation policies, and scheduling strategies, e.g.. see [10, 6, 4, 7], and the references therein. In all these works, the sampling process does not change the dynamics of the underlying Markov chain. Thus, it is important to note that the aforementioned studies do not consider the presence of an external control that can actively influence the state evolution of the system. However, in edge computing scenarios, each job offloading event may alter the state of the device, thereby introducing a control-dependent dynamic. Motivated by this, in the literature, recent studies investigate the problem of tracking a Markov process under external control inputs [1, 9, 2, 8, 12, 3]. These studies are mainly done in the context of job offloading to computing devices whose dynamics can be modeled as a Markov process. Consequently, the term Markov machine has been introduced in the literature to describe such devices.
In this work, we consider a job-assignment system, where $N$ users present in the system stochastically submit their job requests to a central server, as shown in Fig. 1111The icons are taken from www.flaticon.com.. For each user, the central server has a job queue, and it stores the job requests of a user in its dedicated job queue. The system includes a Markov machine that serves a job request assigned by the central server. At the beginning of a time slot, the server may sample the state information of the machine, paying a sampling cost, and if it observes that the machine is in a free state, it schedules a user and submits a job from the corresponding job queue to the machine. As we are considering a job-offloading system, our main goal is to maximize the number of jobs completed per unit time, and to this end, we introduce a novel metric termed the age of job completion. As will be discussed later, neither this metric nor a similar metric has been considered in the existing literature on job assignment to Markov machines. To minimize the age of job completion and the sampling cost, we propose two policy pairs, each comprising a scheduling policy and a sampling policy. For both policy pairs, we numerically find their performance. Furthermore, we derive sufficient conditions that ensure the stability of the job queues for both policy pairs. To the best of our knowledge, such stability results have not been explored in prior studies on job offloading to Markov machines.
@@ II Related Works
In [1], the authors considered a Markov machine equipped with a sampler that samples and transmits its state information to a central server at the cost of sampling. To quantify the tracking performance, the authors introduced a tailored version of the age of incorrect information (AoII) metric. The AoII metric does not directly address the maximization of the number of completed jobs, whereas the metric proposed here, i.e., the age of job completion metric, better captures this objective in job-offloading systems.
@@
In [9], the authors considered a network of Markov machines, where the jobs arrive at a resource allocator following a Poisson process. The authors studied an optimal sampling rate allocation problem for query-based sampling of the machines. To monitor the tracking performance, the authors used the well-known binary freshness (BF) metric. They also introduced the false rejection ratio (FRR) and the false acceptance ratio (FAR) metrics to quantify the effects of limited tracking: the former measures the fraction of jobs rejected despite the presence of a free machine due to insufficient tracking, while the latter measures the fraction of jobs accepted even though no machine was actually free, again due to limited tracking. It is important to note that, similar to AoII, none of these metrics directly captures the objective of maximizing the number of completed jobs. Moreover, the authors in [9] did not consider a job queue at the resource allocator. In [12], the authors considered a job assignment system with a network of exhausted workers and a resource allocator, where the efficiency of each worker depends on its current state. Similar to [9], the authors formulated and studied an optimal sampling rate allocation problem for query-based sampling of these workers, aiming to maximize the probability that each worker operates in its most efficient state.
@@
A revenue maximization problem has been considered in a job assignment system in [8]. They considered that a resource allocator has a unit-size buffer, i.e., it can store at most one job at a time. In [3], the authors considered a job-dispatching system with a network of dispatchers and edge servers; these servers were modeled as Markov machines. Their goal was to maximize the differential gain between job completion and sampling cost, and they formulated the problem as a discounted Markov decision process (MDP). As a result, their work did not directly maximize the average number of jobs completed over the time horizon. When a job was dispatched to a server, it was stored in a limited-size job queue, and if the queue was full, the oldest job was replaced. Similar to [3], the authors of [2] considered a job-dispatching system for a network of dispatchers and edge servers; however, in contrast to [3], they did not consider a job queue at the edge server, i.e., if a dispatcher offloads a job to an edge server and if it is busy, the job is dropped. They proposed an index-based policy that maximizes the probability that a dispatched job is accepted by an edge server. However, in our work, the central server first samples the state of the machine and submits a job only if the machine is free, resulting in an acceptance probability of $1$ for offloaded jobs. Thus, our focus is on maximizing the number of jobs completed per unit time, rather than the job acceptance probability as in [2].
@@
In contrast to [1, 9, 2, 8, 12, 3], in this work, we consider limitless full job queues, making the study of queue stability a necessary aspect of our problem.
@@ III System Model and Problem Formulation
In our job assignment system, user $i$ submits a job request to the central server at the end of each time slot, according to an i.i.d. Bernoulli random variable $a_{i}(t)$, with probability of success being $p_{i}>0$. When the machine is not executing a job assigned by the server, denoted as an external job, the machine can be either free or busy processing an internal job. We model the dynamics between these free and internally busy states with a binary symmetric Markov chain, with probability of state transition being q>0q>0.222For simplicity of computations, in this work, we consider a symmetric state transition, and the general case will be considered in a journal version.
@@
The machine only accepts an external job when it is in the free state. Thus, when the machine is not executing an external job and the server intends to assign a job to the machine, the server samples the state information of the machine, which incurs a cost $L$, and receives the sampled state information instantaneously. We denote this sampling operation at time $t$ with an indicator variable $\mu(t)$, where $\mu(t)=1$ implies that the server samples the state information of the machine at time $t$, and $\mu(t)=0$ implies otherwise. If the server infers that the machine is in a free state, the server instantaneously assigns a job to the machine from one of the non-empty job queues. At time $t$, we denote this scheduling operation with a vector $(\pi_{i}(t))_{i=1}^{N}$. At time $t$, if the server schedules a job from the job queue of user $i$, then $\pi_{i}(t)=1$; otherwise $\pi_{i}(t)=0$. Thus, we represent an employable policy by the central server as
$\phi=(\mu(t),(\pi_{i}(t))_{i=1}^{N})_{t=1}^{\infty}.$ (1)
@@
We assume that the server performs the sampling, scheduling and job assignment processes, at the beginning of a time slot and the machine completes either an internal job or an external job at the end of a time slot. When the machine is processing a job of user $i$, the service time follows a geometric distribution with parameter $q_{i}>0$. During this period, the machine cannot process any other job, and the server is aware of the state of the machine. Once the job is completed, the server is instantaneously notified. After the machine completes a job, it transits to the internally busy state with probability $s>0$, or to the free state with probability (1s)(1-s). Thus, if an external job is completed at time slot t1t-1, the server becomes uncertain about the state of the machine at the beginning of slot $t$ and must sample its state information before submitting a future job.
At time $t$, we denote the last known state information of the machine at the server with $x(t)\in\{-1,0,\ldots,N,f_{r}\}$. Here 1-1 corresponds to the internally busy state, the state 0 denotes the ambiguity of the state of the machine after it completes a job, $i\in\{1,\ldots,N\}$ indicates that the machine is processing a job of user $i$, and $f_{r}$ denotes that the machine is free. Note that, $x(t)$ gets updated if the server samples the state of the machine at time $t$, or receives a job completion acknowledgment at t1t-1, otherwise $x(t)$ evolves as x(t)=x(t1)x(t)=x(t-1). Note that state $f_{r}$ is only achieved when the sampled state of the machine is free and the server does not schedule any job. Furthermore, $x(t)$ remains 0 from the time a job is completed until the server next samples the state information. Similarly, with $x_{a}(t)$, we denote the actual state of the machine at time $t$. The variable $x_{a}(t)$ takes the values $\{-1,1,2\cdots,N,f_{r}\}$, retaining the same interpretation as $x(t)$. Fig. 2 shows the dynamics of the considered Markov machine.
@@
Intuitively, a policy $\phi$ that minimizes (4) also maximizes the number of completed jobs per unit time. This argument is formally shown in a subsequent work. In this work, first we present an adaptive randomized policy that comprises an adaptive randomized sampling policy and an adaptive randomized scheduling policy. Then, we study the max-age scheduling policy together with an adaptive randomized sampling policy.
At time $t$, we observe the non-empty job queues and schedule the corresponding users with a stationary randomized policy. Whenever an identical set of job queues becomes non-empty, the same stationary randomized policy is applied. We call this policy an adaptive randomized scheduling policy. Similarly, based on the non-empty job queues, we define an adaptive randomized sampling policy. We provide a formal definition of an adaptive randomized policy in Definition 1.
@@
An adaptive stationary randomized policy can be viewed as a collection of pairs of stationary randomized policies, one governing user selection and one governing sampling, each supported on a distinct subset of $[N]$.
In a conventional stationary randomized scheduling policy, the central server schedules a user with a fixed probability throughout the entire time horizon, e.g., see [5]. However, at a given time slot, a job queue can be empty; thus, scheduling that user during that slot with a non-zero probability would waste system resources. Thus, using system resources efficiently while being mathematically tractable motivates the proposed adaptive randomized policy. Similar to the definition of an adaptive randomized policy in Definition 1, we formally define an adaptive randomized sampling policy.
@@
$\bar{\Pi}=\{\,\mu(\mathcal{S}):\mathcal{S}\subseteq[N],\,\mathcal{S}\neq\varnothing\,\},$ (9)
@@
At time $t$, the state of the system is described by the vector $\bm{y}(t)=\left((Q_{i}(t))_{i=1}^{N},(v_{i}(t))_{i=1}^{N},\>x(t),\>\gamma(t)\right)$, where $\gamma(t)$ represents the number of slots since the server last knew the true state of the machine. We denote the set of all possible states by $\mathcal{Y}$. Thus,
$\mathcal{Y}=\mathbb{N}^{N}\times\mathbb{N}^{N}\times\{-1,0,\cdots,N,f_{r}\}\times\mathbb{N}.$ (10)
Note that under the proposed policies, $\bm{y}(t)$ evolves as a Markov chain. Next, we define the notion of stability [13].
@@
The system is stable if all recurrent states of the Markov chain $(\bm{y}(t))_{t=1}^{\infty}$ are positive recurrent and the process hits the recurrent states with probability one. Furthermore, when the whole system is stable we say that the queues are also stable.
@@ IV-C Stabilizing the Queues
We now derive sufficient conditions for queue stability under both an adaptive randomized policy and a max-age policy. In both cases, the state of the system $\bm{y}(t)$ evolves as a Markov chain and we follow Definition 3 for stability.
@@
stabilizes the system if there exists $\epsilon>0$ such that, for every non-empty subset $\mathcal{S}\subseteq[N]$,
@@
Note that $1-\chi(q,s)$ provides a lower bound on the probability that the machine is free when sampled. With this interpretation, the conditions in Proposition 1 ensure that the total arrival rate does not exceed the worst-case effective service rate under any of the stationary randomized policies. Conditions of this nature are, in general, known to be necessary and sufficient for rate stability [11]. Here, we further show that they stabilize the queues according to the notion of stability in Definition 3, under the nontrivial dynamics of the considered model.
@@
Under any adaptive randomized policy, the system is stable if there exists $\epsilon>0$, such that
@@
Consider a system with $N$ users having arrival rates $p_{i}$, service rates $q_{i}$, and machine parameters $q$ and $s$. A max-age policy $\bar{\phi}=\left(\mu,\pi^{MA}\right)$ stabilizes the system if there exists $\epsilon>0$ such that for every non-empty subset $\mathcal{S}\subseteq[N]$,
@@ V Numerical Results
We observe numerically that queue stability is achieved only when the users’ arrival rates are well below their respective service rates. For example, with model parameters $N=4$, $q=0.35$, $s=0.3$, $L=5$, service rates $\bar{\bm{q}}=[0.55\;0.73\;0.84\;0.91]$, and arrival rates $\bm{p}=[0.09\;0.09\;0.12\;0.14]$, the conditions in Propositions 1 and 2 are not satisfied and the queues are found numerically to not be stable. Furthermore, the conditions stated in Propositions 1 and 2 are observed to be only sufficient and not necessary for queue stability. For model parameters $N=4$, $q=0.5$, $s=0.5$, $L=5$, service rates $\bar{\bm{q}}=[0.4\;0.6\;0.8\;0.94]$ and arrival rates $\bm{p}=[0.04\;0.05\;0.06\;0.06]$, the sufficient stability conditions are not satisfied; however, numerical results show that the queues remain stable both under $\phi_{1}$ and $\bar{\phi}_{1}$.